Each floor of a square building has an area of 6.4×105 square feet.
What is the length of each side of the floor?
feet
If the building has 12 stories, each with the same floor area, what is the total area of all of the floors?
square feet

Answer by Alan3354(69443)

Each floor of a square building has an area of 6.4×105 square feet.
What is the length of each side of the floor?
-----------
64*105 = 6720 sq ft = s^2
s =~ 81.98 feet
===========================
If the building has 12 stories, each with the same floor area, what is the total area of all of the floors?
6720*12 = 80640 square feet